OpenStudy (anonymous):

integrate arcsin[x] / sqrt[1+x]

OpenStudy (dumbcow):

integration by parts u = arcsin dv = 1/sqrt(1+x) du = 1/sqrt(1-x^2) v = 2sqrt(1+x) \[\int\limits u dv = u v - \int\limits v du\] \[= 2 \sin^{-1} (x) \sqrt{1+x} - 2 \int\limits \frac{dx}{\sqrt{1-x}}\] ...

OpenStudy (anonymous):

thank you, i had my parts flipped to what you have; which complicated it more then simplify :)

OpenStudy (anonymous):

Just to verify your answer \[ \int \frac{\sin ^{-1}(x)}{\sqrt{x+1}} \, dx=\frac{2 \left(2 \sqrt{1-x^2}+(x+1) \sin ^{-1}(x)\right)}{\sqrt{x+1}} + C \]